Daniel had some toys. He gave 1/4 of his toys to Jed, 1/3 to Tim and shared the remainder equally with his other 2 friends. What fraction of his toys did he have left?
Found 2 solutions by greenestamps, Edwin McCravy:
Answer by greenestamps(13331)



0

The statement of the problem says he gave them all away....


Answer by Edwin McCravy(20077)

What's confusing is that the problem as stated doesn't make it clear whether
the 1/3 he gave to Tim was 1/3 of what he had in the beginning, or 1/3 of
what he had left after he gave the 1/4 to Jed.

If it's the first way he gave Jed and Tim 1/4+1/3 = 3/12+4/12 = 7/12 and had
5/12 left. He then shared that 5/12 three ways, 1/3 of 5/12 is 5/36.  So he
gave 5/36 to the first of the other two friends, also 5/36 to the second of
the other two friends, and kept 5/36 for himself.  So this way the answer
would be 5/36.

If it's the second way he gave Jed 1/4 and had 3/4 left.  Then he gave Tim
1/3 of that 3/4.  1/3 of 3/4 is 1/4, which is what he gave to Jed, and had
the other 3/4 of that 3/4 left, and 3/4 of 3/4 is 9/16. So he had 9/16 of
what he had in the beginning left.  He then shared that 9/16 three ways, 1/3
of 9/16 is 3/16.  So he gave 3/16 to the first of the other two friends,
also 3/16 to the second of the other two friends, and kept 3/16 for himself.
So this way the answer would be 3/16.

You should point this out to your teacher, to make sure the wording is not
ambiguous in the future.

Edwin